A force F=(3 i + 4 J + 5 k) newtons is applied on a 500 g object to move it a displacement D=(5 i+ 4 j + 3k) meters, find the angle between the force F and the displacement d, if the work done by f is 46J?
1
Expert's answer
2020-08-05T16:41:48-0400
"W=\vec{F}\cdot \vec{D}=FD\cos(\theta)"

Thus,

"F=\sqrt{3^2+4^2+5^2}=5\sqrt{2}\\
D=\sqrt{3^2+4^2+5^2}=5\sqrt{2}"

Hence,

"46=50\cos(\theta)\implies \theta=\cos^{-1}(23/25)"